Logiweb dictionary of propcalc

Up Help

0 0 propcalc
1 0 nani sete nika...:1

2 0 A1 3 0 A2 4 0 A3

5 0 nani sete nika...:5
6 0 nani sete nika...:6

7 0 MP

8 0 nani sete nika...:8
9 0 mendelson lemma one eight
10 0 nani sete nika...:10
11 0 nani sete nika...:11
12 0 nani sete nika...:12
13 0 nani sete nika...:13

14 0 Mendelson1.8 15 2 ∗ ⇒ ∗

The pyk compiler, version 0.grue.20050603 by Klaus Grue GRD-2005-06-16.UTC:21:00:55.684114 = MJD-53537.TAI:21:01:27.684114 = LGT-4625672487684114e-6